In an experiment, 13.88 g of potassium propionate, C2H5CO2K, were weighed in a beaker and
dissolved in a small amount of distilled water. The mixture is transferred to a 100 mL volumetric
flask, made up to the mark and shaken well. 50.0 mL of this standard solution is mixed with 50.0
mL of a standard solution of propionic acid, C2H5CO2H in order to prepare a buffer solution of
pH 5.66. The molar mass of potassium propionate is 112 g mol-1
and the pKa of propionic acid is
4.87.
(i) What is the concentration of the propionic acid used to make the buffer?
1
Expert's answer
2020-06-08T15:31:09-0400

Concentration of "C_2H_5COOK" in "100" ml "=\frac{13.88}{112\times 100}" "mol\/l"

Now "50" ml of this solution is taken and then mixed with equal amount of propionic acid"(" say "HA)"

So,total volume of buffer solution will be "100 \ ml" .

Concentration of "C_2H_5COOK" or "C_2H_5COO^-" "(" Lets say "A^-)" "=\frac{13.88}{112\times 100}\times \frac{50}{100}=6.2\times 10^{-4}\ mol/l"

Now ,for buffer solution,

"pH=5.66;pK_a=4.87" and

"pH=pK_a+log_e\frac{[A^-]}{[HA]}\implies" "5.66=4.87+log_e\frac{6.2\times 10^{-4}}{[HA]}"

"\implies" "log_e\frac{6.2\times 10^{-4}}{[HA]}=0.79\implies\frac{6.2\times 10^{-4}}{[HA]}=e^{0.79}=2.2034"

"\implies [HA]=\frac{6.2\times 10^{-4}}{2.2034}=2.81\times 10^{-4}\ mol/l"